Target meeting: 20May09, 09:00

Present: CB, RN, PH, PS, JT, KL, JL, AA

1. Review of actions:

·         JT/RN: Clarify issue of clearance etc. raised above via drawing and email.
JT had sent drawing to RN.  2.4—2.5mm clearance; need to check tolerances at assembly.  Optical housing fit will need to be dealt with at fit-up.  Believed to be possible to solve the problem by modifying or changing a screw head.  OK to carry on.

·         JT/PS: Agree final specification of marker detail.
Done.

·         JT: Modify assembly drawing to show correct orientation of vane.
Done.

·         JT/RN: Clarify clearance required by connector blocks and ensure design of cover is commensurate with the required clearance.
Drawing of assembly sent to RN.  Only affects electrical safety covers, which will be retrofitted.  So, believe OK to carry on.

·         PH/PS: Meet with S.Griffiths et al at DL to discuss details of electrical integration in R78 and ISIS.
Meeting took place yesterday.  Make agenda item on electrical.

·         JT: Seek to book high-speed camera for relevant period.
Stands.

·         CB: Collate team’s holiday plans.
Done.

2. Progress on mechanical design and fabrication:

2.1 Target assembly: JT

·         AC: core parts are being machined – at point of making final cuts;

·         RH: making bottom flange;

·         Tallirond has been calibrated and used on test pieces;

·         Bearing surface metrology has been tested – final adjustment to height required;

·         Parts for welding dropped off at Culham.  Contract not sent (%&*?!!); contractual issue has been addressed.

·         Vacuum tubes brazed and sent to Sheffield (they have arrived);

·         Stop weld and associated document complete;

·         Have received veins.  Have 40, first one out of packet was of poor quality.  Rest OK, three from the ‘best batch’ sent to Sheffield for test on target rig;

·         All shaft drawings now released;

·         Grinding order placed yesterday;

·         Principal worry today is that delay of order to Culham may have delayed welding.  JT will check today;

·         Material for 2-piece shaft has still not arrived (%&*?!!).  JT will chase.

2.2 Stator: CB

·         Stator has been awaiting vacuum tube.  Tube now there so next job is to transfer coils on to tube;

·         Resistance check has been carried out and was OK;

·         Will fit tube in the next couple of days.  Also require retaining cylinders.  JT has sourced the cylinders and will chase to make sure they get to Sheffield next week;

·         Still on schedule to supply two stators for CM24.

2.3 Modifications to gate valve and vacuum vessel: EMcC

·         No report this week.

3. Progress on electrical installation etc.

·         S.Griffiths will organize power electronics and electrical upgrade in June.  Will do equipment in R78 in week beginning 22Jun and equipment in ISIS in week of 29 Jun;

·         Equipment will be identical so that we have a set of active spares in R78;

·         Some drawings need to be updated to be completed by the end of June;

·         Thermocouple (K-type) spare channel is available that terminates in the vac-rack area;

·         Requested labeling and documentation of terminations in ISIS to be done.  Will check and make inventory at the beginning of the collaboration meeting.  CB will coordinate with EMcC and JL;

·         Monitoring of chiller: PH to ask WS of present status;

·         Long fibre-optic cables: presently at risk because have three runs of cables.  Spare fibre-optic cables were laid.  On one route there are no spare cables.  Therefore need to pull some additional cables as spare;

·         Actions:

o   CB: Coordinate with EMcC and JL labeling and cable documentation.

o   PS: Make proposal for fibre spares and get a quote;

o   PH Plan schedule/plan for spare fibre laying

3. Progress on (dynamic) mechanical FEA: SY, WL

·         No report this week.

4. Progress on magnetic FEA: PS

·         No report this week.

5. Target electronics and DAQ: PS, JL

Controller/USBDAQ interface boards:

·         Progress on board manufacture, PS will begin work on VHDL coding ~today.

6. Commissioning plan: discussion

·         Updated version in preparation.  Holidays and details of schedule being taken into account;

7. Next meeting

·         Next target meeting: 27May09

·         Next target w/s, with ISIS: possible dates in June 17 (preferred), 18

1.    ISIS contributions:

1.    Interlocks/use of signals on ISIS side;

2.    Beam loss monitor recording within ISIS;

8. AoB

·         PH: Question about network; check with Craig MacWaters.

Summary of actions:

·         CB: Coordinate with EMcC and JL labeling and cable documentation.

·         PS: Make proposal for fibre spares and get a quote;

·         PH: Plan schedule/plan for spare fibre laying

·         JT: Seek to book high-speed camera for relevant period